# Learning noisy tissue dynamics across time scales

### v1

@@ Abstract
Predicting dynamic cell signaling. Finally we wanted to test if the GNN could be used to model the propagation of cell-cell signaling across a tissue. In these signaling systems, both the deterministic drift [Uncaptioned image] and the noise [Uncaptioned image] in Eq. 1 are equally important. The activation of a cell causes a signal to be passed to neighbor cells which activates them and in turn propagates the signal forward. Predicting such processes requires knowledge of the history of the system (represented by a stack of frames in Fig. 3A). In this case, the combination of GNN with WaveNet (GWN) is necessary because conventional GNNs fail. At each layer of the model, a GNN is used to encode the spatial dependencies between nodes in the graph, while the dilated convolutional layers capture the temporal dependencies between samples.
The combination of these two types of layers allows the model to predict the propagation of ERK (extracellular signal-regulated kinase) signaling, a signaling pathway involved in the regulation of cell division in differentiated cells Aoki et al. (2013); Hino et al. (2020); Ram et al. (2023), see Fig. 3A. Recently, ERK activity sensors have been developed and it was discovered that ERK signaling produces mechanochemical waves that propagate through a tissue Boocock et al. (2020); Hino et al. (2020). However, models of this process are limited to continuum like approaches Boocock et al. (2020) due to the challenge of simultaneously predicting the stochastic ERK signals and the motion of cells which occur at the same timescale.
@@
We use a previously published dataset of ERK dynamics in epithelial MDCK monolayers and segment the cells to track the position, nuclear morphology and ERK signaling state of each cell Boocock et al. (2020). We trained the GWN on a time series of this data, and then provided with an unseen test set. The first 7 frames of the test set are provided to the GWN because history is required to make further predictions. Next, we query the model to predict the future signaling dynamics on unseen samples. We find that the GWN produces an output of wavelike dynamics that mimic the experiment (Fig. 3B-C). As the propagation of these waves is highly sensitive to the initial conditions and noise in the system, a bit like turbulent waves in the sea, the exact wave pattern cannot be predicted out to long time scales. Nonetheless, the predicted signaling patterns match the experiment in magnitude and fluctuations (Fig. 3D). Further, when we analyze the spatial and temporal correlations of the signaling waves we see that both the magnitude and variation in these correlations match the experiment (Fig. 3E-F). This suggests that the GWN has learned the same rules of signal propagation that are present in the experiment.
@@
In this example, our algorithms reveal the different time scales at play in the tissue. Figure 3D shows evolution on a long time scale of the order of hundreds of minutes. In this case, the network captures long-term changes in the system, likely occurring through accumulated cell division and rearrangement that alter the structure of the tissue. On the other hand, Fig. 3E shows how the network captures the two shorter time scales associated with the oscillation and decorrelation of the waves, of the order of 10 to 20 minutes. These timescales are associated with protein signaling dynamics that occur within the cell and are passed to neighboring cells. In short, our generative model learns the mechanochemical coupling between active stress and the ERK pathway that underlie the spatiotemporal patterns in the MDCK monolayers and gives access to all many-body correlations.
Outlook. To sum up, we have developed a framework to learn neural stochastic differential equations describing the dynamics of interacting cells in biological tissues. Our method, based on graph neural networks, is geared towards discrete data, not restricted to biology, such as individual cells or particles with unknown interactions. The resulting description encompasses both physical and biological degrees of freedom, including cell states and signaling pathways.
@@
Our work paves the way towards data-driven methods that could disentangle the complex multicellular signaling networks at play in tissues. One of the current major challenges in biology is to measure the gene expression and regulatory landscape of individual cells, and to integrate this information with spatial data describing the environment of the cell within tissues and organs Bressan et al. (2023); Levsky et al. (2002); Satija et al. (2015); Wolf et al. (2018); Kulkarni et al. (2019); Haviv et al. (2024). The efficient inference of single-cell properties of our algorithm combined with existing dimensionality-reduction methods for single-cell gene expression data could directly yield the reduced representation of cells states and transitions within tissues. By describing the tissue as a whole through a joint probability distribution, rather than starting from the gene expressions of isolated cells, this would allow one to capture correlations between different cells from the get-go.
@@
Furthermore, our trained neural networks could be used as a digital twin that facilitates the diagnosis of diseases such as tissue inflammation. This would entail first calibration with lab data, where many biological markers are measured, and subsequently deployment on patient data, where a more limited amount of measurements is available.

### v2

@@ Abstract
Dynamic cell signaling in ERK waves. Finally we deploy our algorithms in situations where the propagation of cell signaling across a tissue is stochastic, a bit like turbulent waves at sea. In these cases, both the deterministic drift [Uncaptioned image] and the noise [Uncaptioned image] in Eq. 1 are equally important. The activation of a cell causes a signal to be passed to neighbor cells which activates them and in turn propagates the signal forward. Predicting such processes requires knowledge of the history of the system (represented by a stack of frames in Fig. 3A). In this case, the combination of GNN with WaveNet (GWN) is necessary because conventional GNNs fail. At each layer of the neural network, a GNN is used to encode the spatial dependencies between nodes in the graph, while the dilated convolutional layers capture the temporal dependencies between samples. The combination of these two types of neural network layers allows the model to predict the propagation of ERK (extracellular signal-regulated kinase) signaling, a signaling pathway involved in the regulation of cell division in differentiated cells Aoki et al. (2013); Hino et al. (2020); Ram et al. (2023), see Fig. 3A. Recently, ERK activity sensors have been developed and it was discovered that ERK signaling produces mechanochemical waves that propagate through a tissue Boocock et al. (2020); Hino et al. (2020); Boocock et al. (2023). However, models of this process are limited by the challenge of simultaneously predicting the stochastic ERK signals and the motion of cells, which occur at the same timescale. We are not aware of models that can quantitatively reproduce both the average evolution and the stochastic correlations, nor inferring these quantities from data in a model agnostic manner. Here we ask: can we bridge this gap with our dual graph WaveNet framework.
@@
We use a previously published dataset of ERK dynamics in epithelial MDCK monolayers and segment the cells to track the position, nuclear morphology and ERK signaling state of each cell Boocock et al. (2020). We trained the GWN on a time series of this data, and then provided it with an unseen test set. The first 7 frames of the test set are provided to the GWN because history is required to make further predictions. Next, we query the model to predict the future signaling dynamics on unseen samples.
We find that the GWN produces an output of wavelike dynamics that mimic the experiment (Fig. 3B-C). As the propagation of these waves is highly sensitive to the initial conditions and noise in the system, the exact wave pattern cannot be predicted out to long time scales. Nonetheless, the predicted signaling patterns match the experiment in magnitude and fluctuations (Fig. 3D). Further, when we analyze the spatial and temporal correlations of the signaling waves we see that both the magnitude and variation in these correlations match the experiment (Fig. 3E-F). This suggests that the same variables (cell morphology and positions plus ERK activity) that are relevant to the deterministic dynamics modeled in Ref. Boocock et al. (2020) (i.e., that allow one to predict the future of the system) control the stochastic dynamics of the system.
@@
Beyond this qualitative conclusion, our algorithm reveals the different time scales at play in the tissue: Fig. 3D shows evolution on a long time scale of the order of hundreds of minutes while Fig. 3E shows how the network captures the two shorter time scales associated with the oscillation and decorrelation of the waves, of the order of 10 to 20 minutes. It is the interplay between the protein signaling dynamics on short time scales and the accumulation of many discrete events (cell division and rearrangement) that make the tissue move on the longer time scales. To sum up, our generative model discovers directly from data the mechanochemical coupling between active stress and the ERK pathway that underlie the spatiotemporal patterns in the MDCK monolayers and gives access to all many-body correlations.
@@
Outlook. One of the current major challenges in biology is to measure gene expressions and regulatory landscapes of individual cells, and to integrate this information with spatial data describing the environment of the cell within tissues and organs Bressan et al. (2023); Levsky et al. (2002); Satija et al. (2015); Wolf et al. (2018); Kulkarni et al. (2019); Haviv et al. (2024). Our work paves the way towards a holistic description of tissues, through a joint probability distribution, rather than starting from the gene expressions of isolated cells. By integrating this multicellular context with single-cell gene expression data, our approach could directly yield the reduced representation of cell state transitions in a way that captures multiscale correlations from the get-go.
@@
Furthermore, our trained neural networks could be used as a digital twin in clinical studies of tissue inflammation. This would entail first calibration with lab data, where many biological markers are measured, and subsequently deployment on patient data, even if only a more limited amount of measurements is available.